# Ola: Pushing the Frontiers of Omni-Modal Language Model with Progressive Modality Alignment

### v2

@@ Abstract
Recent advances in large language models, particularly following GPT-4o, have sparked increasing interest in developing omni-modal models capable of understanding more modalities. While some open-source alternatives have emerged, there is still a notable lag behind specialized single-modality models in performance. In this paper, we present Ola, an Omni-modal language model that achieves competitive performance across image, video, and audio understanding compared to specialized counterparts. The core design of Ola lies in its progressive modality alignment strategy that extends the supporting modality of the language model progressively. Our training pipeline begins with the most distinct modalities: image and text, then gradually expands the skill sets of the model using speech data that connects language and audio knowledge, and video data that connects all modalities. The progressive learning pipeline also enables us to maintain a relatively small size of the cross-modal alignment data, making developing omni-modal from existing vision-language models easy and less costly. Moreover, to unlock an advanced interactive experience like GPT-4o, we further design a sentence-wise decoding solution for streaming speech generation. Extensive experiments demonstrate that Ola surpasses existing open omni-modal LLMs across all modalities while achieving highly competitive performance compared to state-of-the-art specialized models of similar sizes. We aim to make Ola a fully open omni-modal understanding solution to advance future research in this emerging field. Model weights, code, and data are open-sourced at https://github.com/Ola-Omni/Ola.
Figure 1: Ola pushes the frontiers of the omni-modal language model across image, video and audio understanding benchmarks. We compare Ola with existing state-of-the-art open-sourced multimodal models and GPT-4o on their abilities in mainstream image, video, and audio benchmarks. For fair comparisons, we select around 7B versions of existing MLLMs. Ola can achieve outperforming performance against omni-modal and specialized MLLMs in all modalities thanks to our progressive alignment strategy. “$\times$” indicates that the model is not capable of the task and “--” indicates the result is lacking. The score for LibriSpeech is inverted as lower is better for the WER metric.
@@ 1 Introduction
Figure 2: Ola Architecture. Ola supports omni-modal inputs including text, image, video, and audio, capable of processing the inputs simultaneously with competitive performance on understanding tasks for all these modalities. Meanwhile, Ola supports user-friendly real-time streaming decoding for texts and speeches thanks to the text detokenizer and the speech decoder.
@@
Figure 3: Progressive modality alignment helps to learn better omni-modal models. We compare our progressive alignment strategy with two baseline training pipelines on Image QA(MMBench [40]), Video QA(VideoMME [21]), and ASR(LibriSpeech [54]): 1) direct mixing where all instruction tuning data is merged and trained in a single stage, and 2) balanced sampling where we upsample certain sources to make the training data more balanced among modalities. The experiment is conducted on a subsampled training set for efficiency and we train models for the same number of steps for fair comparisons. The score is normalized based on the score of progressive alignment to calculate the relative score and the ASR score is inverted as lower is better for the WER metric.
@@
The core challenges in training omni-modal Large Language Models lie in the modeling of modalities in various distributions, and the design of an effective learning pipeline to achieve competitive, balanced performance on all the supported tasks. Several attempts have been made to overcome the difficulty of omni-modal models [22, 17, 72], where we illustrate the mainstream works and the state-of-the-art MLLMs [67, 84, 32, 11] in specific domains in Fig. 1. Impressive performance and modality breadth are contradicted in previous works, while existing open-sourced omni-modal solutions still have a large performance gap between state-of-the-art specialized LLMs, making a strong barrier between the concept of omni-modal and real-world applications. Moreover, the lack of capability in specific domains or tasks, the mass data demand, the delay for user interaction, and the inadequate alignment across modalities show the suboptimal for existing omni-modal models.
In this paper, we propose the Ola model, exploring the solution for training an omni-modal Large Language Model with comparable performance with state-of-the-art specific LLMs, real-time interaction, and high efficiency on alignment data. The core design of the Ola model is the progressive modality alignment strategy. To build connections between language and vision, we start from two fundamental and separated modalities, image and text, to build the basic knowledge for omni-modal models. Subsequently, we gradually expand the training sets to equip the model with an extended ability including video frames that strengthen the visual understanding capability, speech data that connects the language and audio knowledge, and the video with audio that mixes up the information from language, video, and audio comprehensively. The progressively learning strategy makes omni-modal learning easier by disassembling the complex training procedure into small steps, therefore maintaining a small size of cross-modal alignment data and making it easier to start from existing achievements in vision-language models. As shown in Fig. 3, the performance of Ola largely benefits from our progressive training pipeline, leading to more balanced and competitive results on all modalities.
To cooperate with the training strategy, important improvements have been made to the architecture and data domains. 1) The Ola architecture supports omni-modal inputs and streaming text and speech generation with extendable and concise architecture design. We design the joint alignment module for vision and audio, fusing the visual inputs with a local-global attention pooling layer, and make free combinations for visual, audio, and text tokens. Moreover, we integrate the sentence-wise streaming decoding module for high-quality voice synthesis. 2) Beyond the collected fine-tuning data in vision and audio aspects, we deeply excavate the relationships between video and the corresponding audio to construct the bridge between visual and audio modality. Specifically, we collect raw videos from academic and open-ended web sources, design separated cleaning pipelines, and then utilize vision-language models to generate question-answering pairs based on the subtitles and video content.
@@
We evaluate Ola under the complete omni-modal benchmarks including image, video, and audio aspects. With only 7B parameters, Ola achieves competitive performance across mainstream multi-modal benchmarks. On Image Benchmarks, Ola excels at general understanding and specific-task understanding, with an overall mean accuracy of 72.6% on the challenging OpenCompass benchmark [16], 84.3% average scores on MMBench-1.1 [40], 57.0% average scores on MMMU [77], etc. On the challenging VideoMME [21] multiple-choice benchmark ranging from videos within 30 seconds to 1 hour, Ola achieves the impressive accuracy of 68.4% with video and audio inputs. Ola also excels at audio understanding tasks such as audio-speech recognition and chat evaluation, achieving 3.1 mean WER on LibriSpeech [54] and 6.41 GPT-eval score on AIR-Bench [74]. Results on the benchmarks show a giant promotion compared with existing omni-modal LLMs and even outperforming the performance of state-of-the-art specialized LLMs.
@@ 2 Related Works
Large Vision-Language Models. Inspired by the success of AI assistants and large language models [50, 51, 52], research has increasingly focused on vision-language multi-modal large language models. Significant advancements have been made in architecture design [3, 33, 39, 13, 44], training strategies [10, 43], model scaling [32], and data curation [31, 45, 43, 73]. Furthermore, models are evolving beyond static images to support video [36, 47, 7, 42], 3D [27, 37], and mixed visual inputs [59, 57]. However, extending these models to effectively integrate audio modalities while maintaining balanced and robust performance remains an area that has not been fully explored.
@@
Towards Large Omni-Modal Models. Recent advancements in large language models [53, 24] have spurred interest in developing omni-modal models that can handle multiple modalities simultaneously. Notable examples include SpeechGPT [80] and LLaMA-Omni [17], which integrate audio-text understanding with speech generation. The Mini-Omni [72] series addresses challenges in speech streaming generation through parallel decoding techniques. VITA [22] extends this capability by unifying audio, image, video, and text understanding. While these models excel in understanding tasks, efforts are also being made to tackle both understanding and generation tasks [71, 68]. However, existing omni-modal models often fall short in managing the full spectrum of input modalities and output formats, or they suffer from significantly poorer performance. Ola aims to address these limitations by enhancing the capability and efficiency of omni-modal models with better architecture, training strategy, and targeting data preparation.
@@ 3 Ola: Omni-Modal Understanding
Figure 4: Illustrations of the Ola Progressive Modality Alignment. We visualize the relationships among modalities in the left part. Speech acts as the connection between language and audio knowledge, while video constructs the bridge with highly relevant visual and audio information. Therefore, we design the progressive alignment training strategy from primary to periphery. Furthermore, we design the cross-modality video-audio data to better capture the relationships among modalities.
We put main efforts into three aspects to obtain an omni-modal understanding for Ola, capable of reading, hearing, seeing, typing, and speaking arbitrarily. 1) The Ola architecture introduced in Sec. 3.1 supports omni-modal inputs and streaming outputs for both text and speech. 2) We design the progressive training strategy in Sec. 3.2 to bridge the modality gaps between language and vision from primary to periphery. 3) Effective omni-modal training data in Sec. 3.3 provide strong performance across all the benchmarks, especially with our proposed cross-modal video data that stresses to learn audio-relevant information from videos.
3.1 Ola Architecture
A general view of Ola architecture is illustrated in Fig. 2. The encoding part accepts omni-modal inputs in text, image, video, and audio formats with modal-specific encoders or embeddings. Subsequently, the joint alignment operations process all the inputs in a unified manner, fusing and concatenating all the sequences into flattened tokens for the core Ola Large Language Model. The LLM generates text tokens in serial, and we adopt the speech decoder for streaming speech decoding.
@@ 3.1 Ola Architecture
Omni-Modal Inputs Encoding. We encode visual, audio, and text inputs separately based on the successful practice of the previous text-to-one-modal Large Language Models. For visual inputs that include images $I$, videos $V_{1,2,\cdots,n}$ with $n$ frames, we follow vision-language models [39, 33] to use multi-modal visual encoder $\mathcal{E}_{I,V}(I,V_{f_{1},f_{2},\cdots})$ for encoding. Note that we preserve the original aspect ratio of each image or frame for the arbitrary resolution vision encoder OryxViT [42] initialized from SigLIP-400M [79], as OryxViT performs a more natural solution for visual inputs. We obtain the image features $f_{I},$ and the video features for each frame $[f_{V_{1}},f_{V_{2}},\cdots,f_{V_{n}}]$ based on the image patches.
For audio encoding, we propose the dual encoder approach for the audio input. Specifically, we use Whisper-v3 [58] as the speech encoder and BEATs [8] as the music encoder for better alignment between audios and texts, providing richer audio information. The music encoder takes the original wav audio $A$ as inputs, while the speech encoder takes the wav transformed into Mel spectrogram representation $A_{(mel)}$ as inputs. Note that the Whisper encoder only supports a certain length of audio inputs, therefore we fix the sample rate as 16000Hz and cut the overlong audio into segments of 30 seconds $A_{1},A_{2},\cdots,A_{n}$ and conduct the encoder operation in batches $[f_{A_{1}},f_{A_{2}},\cdots,f_{A_{n}}]=\mathcal{E}_{A}([A_{1},A_{2},\cdots,A_{n}])$. The embedding features of the speech and music encoders are concatenated across channel dimensions for the comprehensive audio features $f_{A}$.
For the text inputs, we use the carefully designed tokenizer and the well-trained embedding layers from the pre-trained Large Language Model for the text tokens $t_{T}$ directly.
@@
Joint Alignment for Vision and Audio. The alignment modules act as the converter from specific-modal spaces to the text embedding space, which is an essential part of omni-modal Large Language Models. To reduce the token length of visual features for higher efficiency, we obtain one step forward based on the motivation of structural downsampling in previous works [35], and propose the Local-Global Attention Pooling layer for better downsampled features with less information loss. Specifically, for image or video frame feature in spatial shape $H\times W$ and channel $C$, we adopt the bilinear interpolation for 2x downsampling to obtain $f^{\text{global}}$, which contains the global information of the downsampled region. We combine the original and global features for local-global embeddings and use Softmax to predict the importance $\pi$ of each downsampled spatial region:
@@
The downsampled feature $f^{\text{global}}$ determines the weight of each previous region based on the score $\pi$ with the Hadamard product.
@@
Streaming Speech Generation. We adopt CosyVoice [15] as a high-quality speech decoder for speech generation. To support user-friendly streaming decoding, we detect the generated text tokens in real time and truncate the sentence once punctuation is encountered. Afterward, the previous sentence is fed into the speech decoder for audio synthesis. Therefore, Ola does not need to wait for the whole sentence to finish while supporting streaming decoding. Though several attempts [17, 72] have been made to train the speech generation module end-to-end, the external text-to-speech decoder is a more efficient, high-quality, and training-free solution for omni-modal models.
3.2 Progressive Omni-Modal Alignment

### v3

@@ Abstract
Figure 2: Ola Pushes the Frontiers of the Omni-Modal Language Model across Image, Video and Audio Understanding Benchmarks. We compare Ola with existing state-of-the-art open-sourced multimodal models (InternVL-2.5 [9] and Qwen2.5-VL [62]), omni-modal models (Qwen2.5-Omni [69] and VITA-1.5 [22]), and GPT-4o [50]. For fair comparisons, we select around 7B versions of existing MLLMs. Ola can achieve outperforming performance against omni-modal and specialized MLLMs in all modalities. “$\times$” indicates that the model is not capable of the task and “--” indicates the result is lacking. The score for LibriSpeech is inverted as lower is better for the WER metric.
Recent advances in large language models, particularly following GPT-4o, have sparked increasing interest in developing omni-modal models capable of understanding more modalities. While some open-source alternatives have emerged, there is still a notable lag behind specialized single-modality models in performance. In this paper, we present Ola, an Omni-modal Language model that achieves competitive performance across image, video, and audio understanding compared to specialized counterparts, pushing the frontiers of the omni-modal language model to a large extent. We conduct a comprehensive exploration of architectural design, data curation, and training strategies essential for building a robust omni-modal model. Ola incorporates advanced visual understanding and audio recognition capabilities through several critical and effective improvements over mainstream baselines. Moreover, we rethink inter-modal relationships during omni-modal training, emphasizing cross-modal alignment with video as a central bridge, and propose a progressive training pipeline that begins with the most distinct modalities and gradually moves towards closer modality alignment. Extensive experiments demonstrate that Ola surpasses existing open omni-modal LLMs across all modalities while achieving highly competitive performance compared to state-of-the-art specialized models of similar sizes. We aim to make Ola a fully open omni-modal understanding solution to advance future research in this emerging field. Model weights, code, and data are open-sourced at https://github.com/Ola-Omni/Ola.
Figure 1: Roads to Ola Omni-Modal Language Model. We illustrate our innovations on visual and audio understanding in subfigures (a) and (b). Benefiting from the improved architectures and data, Ola achieves comparable performance with specialized models, outperforming state-of-the-art omni-modal models. We carefully design the training strategy for omni-modal models based on cross-modality and progressive alignment, as illustrated in subfigure (c). We compare our approach with two common baselines when training omni-modal models: 1) direct mixing where all instruction tuning data is merged and trained in a single stage, and 2) balanced sampling where we upsample certain sources to make the training data more balanced among modalities.
@@ 1 Introduction
The core challenges in training omni-modal Large Language Models lie in the modeling of modalities in various distributions, and the design of an effective learning pipeline to achieve competitive, balanced performance on all the supported tasks. Several attempts have been made to overcome the difficulty of omni-modal models [21, 16, 68, 69], where we illustrate the mainstream works and the state-of-the-art MLLMs [65, 80, 31, 69] in specific domains in Fig. 2. Impressive performance and modality breadth are contradicted in previous works, while existing open-sourced omni-modal solutions still have a large performance gap between state-of-the-art specialized LLMs, making a strong barrier between the concept of omni-modal and real-world applications. Moreover, the lack of capability in specific domains or tasks, the mass data demand, and the inadequate alignment across modalities show the suboptimal for existing omni-modal understanding models.
In this paper, we propose the Ola model, exploring a comprehensive solution for building an omni-modal Large Language Model with comparable performance to state-of-the-art specific LLMs. We first put efforts into enhancing fundamental comprehension capabilities across multimodal domains, which we address through two key aspects: architecture design and data optimization. 1) The Ola framework features an extendable yet concise architecture supporting omni-modal inputs. Specifically, we develop visual and audio encoders capable of processing diverse inputs, including images, videos, speech, and music. To enable effective cross-modal integration, we design a joint alignment module incorporating local-global attention pooling for visual inputs, while enabling flexible combinations of visual, audio, and text tokens. 2) We systematically curate high-quality multimodal datasets to achieve tier-1 benchmark performance. This includes designing vision-language pre-training strategies that enhance both ViT and LLMs, and we meticulously assemble supervised fine-tuning datasets for both visual and audio modalities.
Based on a strong multi-modal understanding model, we further explore the techniques for combining all the modalities towards a comprehensive omni-modal language model without performance drop. Based on our observation across modalities, we find that video can act as the core for bridging all the modalities, as visual, audio, and video subtitles show a high correlation in one sample. Therefore, we can deeply excavate the relationships between video and the corresponding audio to construct the bridge between visual and audio modality. Specifically, we collect videos from academic and open-ended web sources, design separate cleaning pipelines, and then utilize vision-language models to generate question-answering pairs based on the subtitles and video content. To operate omni-modal training, we design the progressive modality alignment strategy, making omni-modal learning easier by disassembling the complex training procedure into small steps, therefore maintaining a small size of cross-modal alignment data and making it easier to start from existing achievements in vision-language models. As shown in Fig. 1 (c), the performance of Ola largely benefits from our progressive training pipeline, leading to more balanced and competitive results on all modalities.
@@
We evaluate Ola under the complete omni-modal benchmarks, including image, video, and audio aspects. With only 7B parameters, Ola achieves competitive performance across mainstream multi-modal benchmarks. On Image Benchmarks, Ola excels at general understanding and specific-task understanding, with an overall mean accuracy of 72.6% on the challenging OpenCompass benchmark [15], 84.3% average scores on MMBench-1.1 [37], 57.0% average scores on MMMU [73], etc. On the challenging VideoMME [20] multiple-choice benchmark ranging from videos within 30 seconds to 1 hour, Ola achieves the impressive accuracy of 68.4% with video and audio inputs. Ola also excels at audio understanding tasks such as audio-speech recognition and chat evaluation, achieving 3.1 mean WER on LibriSpeech [51] and 6.41 GPT-eval score on AIR-Bench [70]. Results on the benchmarks show a giant promotion compared with existing omni-modal LLMs and even outperforming the performance of state-of-the-art specialized LLMs.
@@ 2 Related Works
Large Vision-Language Models. Inspired by the success of AI assistants and large language models [47, 48, 49], research has increasingly focused on vision-language multi-modal large language models. Significant advancements have been made in architecture design [3, 13, 36, 41], training strategies [10, 40], model scaling [31], and data curation [30, 42]. Furthermore, models are evolving beyond static images to support video [34, 44, 7], 3D [26], and mixed visual inputs [56, 54]. However, extending these models to effectively integrate audio modalities while maintaining balanced and robust performance remains an area that has not been fully explored.
@@
Towards Large Omni-Modal Models. Recent advancements in large language models [50, 23] have spurred interest in developing omni-modal models that can handle multiple modalities simultaneously. Notable examples include SpeechGPT [76] and LLaMA-Omni [16], which integrate audio-text understanding with speech generation. VITA [21, 22] and Qwen2.5-Omni [69] extends this capability by unifying audio, image, video, and text understanding. However, existing omni-modal models often fall short in managing the full spectrum of input modalities and output formats, or they suffer from significantly poorer performance. Ola aims to address these limitations by enhancing the capability and efficiency of omni-modal models with better architecture, training strategy, and data preparation.
@@ 3 Ola: Omni-Modal Understanding
In this section, we present our innovations in building a comprehensive multimodal framework capable of processing arbitrary visual, auditory, and textual inputs. We illustrate the architecture of Ola in Fig. 3. Building upon a pre-trained Large Language Model foundation, we first elaborate our designs for visual content processing in Section 3.1 and audio content integration in Section 3.2. We further conduct an in-depth investigation of modality alignment challenges in omni-modal learning through Section 3.3. To enable effective cross-modal learning, we establish video as the central medium by constructing cross-modal video datasets that emphasize audio-visual correlation learning. Moreover, we design the progressive training strategy to bridge the modality gaps between language and vision from primary to periphery.
Figure 3: Ola Architecture. Ola supports omni-modal inputs including text, image, video, and audio, capable of processing the inputs simultaneously with competitive performance on understanding tasks.
3.1 Advanced Visual Understanding
As illustrated in Fig. 1 (a), starting from a fundamental vision-language model [31], we present a high-quality and state-of-the-art visual understanding model through our improvements on architectures, pre-training, and fine-tuning data.
@@ 3.1 Advanced Visual Understanding
Visual Encoding. For visual inputs that include images $I$, videos $V_{1,2,\cdots,n}$ with $n$ frames, the visual encoding part transforms the pixels into embeddings at language spaces. In the Ola visual encoding, we preserve the original aspect ratio of each image or frame for the arbitrary resolution vision encoder OryxViT [39] initialized from SigLIP-400M [75], as OryxViT performs a more natural solution for visual inputs. With multi-modal visual encoder $\mathcal{E}_{I,V}(I,V_{f_{1},f_{2},\cdots})$ for encoding, we obtain the image features $f_{I},$ and the video features for each frame $[f_{V_{1}},f_{V_{2}},\cdots,f_{V_{n}}]$ based on the image patches.
@@
The alignment modules act as the converter from specific-modal spaces to the text embedding space. Ola treats images and video frames equally to hold the unification of visual contents. To reduce the token length of visual features for higher efficiency, we take one step forward based on the motivation of structural downsampling in previous works [33], and propose the Local-Global Attention Pooling layer for better downsampled features with less information loss. Specifically, for image or video frame feature in spatial shape $H\times W$ and channel $C$, we adopt the bilinear interpolation for 2x downsampling to obtain $f^{\text{global}}$, which contains the global information of the downsampled region. We combine the original and global features for local-global embeddings and use Softmax to predict the importance $\pi$ of each downsampled spatial region:
@@
The downsampled feature $f^{\text{global}}$ determines the weight of each previous region based on the score $\pi$ with the Hadamard product, therefore assigning more weights to the more informative part during the downsampling process.
Visual Pre-Training at Scale. We approach scalable visual pre-training through two key components: ViT pre-training and instruction pre-training. The ViT pre-training is independently conducted on vision transformers, integrated with a small LLM as a language interface. The primary objective of this phase is to enhance the visual-language capabilities of the visual encoder. To achieve this, we utilize large-scale data pairs derived from OCR, grounding, and captioning datasets. The subsequent instruction pre-training stage aims to equip the Ola model with comprehensive knowledge. We collect approximately 20M text-image pairs from both open-source and in-house sources to establish foundational capabilities. To ensure data quality, state-of-the-art vision-language models are employed for re-captioning and re-questioning of the data pairs, ultimately generating a substantial volume of high-quality, instruction-level pre-training data. The detailed data composition and modification pipeline are stated in the Appendix.
Supervised Fine-Tuning. The supervised fine-tuning visual data is collected from academic datasets in image and video categories. For the image data, we follow the simple setting in [35] for image MLP alignment. The MLP alignment data includes 800k image captioning pairs from the LAION dataset [60]. For text-image supervised fine-tuning data, we collect abundant data from various tasks, including captions, conversations, OCR, etc. The source of the training data involves the mixture of LLaVA-OneVision [31], Cauldron [30], Cambrian-1 [65], MAmmoTH-VL [25], PixMo [12], etc., resulting in 7.3M image training data in total.
For text-video training data, we collect useful video datasets from LLaVA-Video-178K [80], VideoChatGPT-Plus [44], LLaVA-Hound [78], and Cinepile [57], with 1.9M video conversation pieces in total. We randomly sample 2/3 video-language data pairs from LLaVA-Video-178K, resulting in 1.2M high-quality training data, and we use the full set of other data sources.
Figure 4: Illustrations of the Relations between Modalities. We visualize the relationships among modalities in the left part. Speech acts as the connection between language and audio knowledge, while video constructs the bridge with highly relevant visual and audio information. Therefore, we design the cross-modality video-audio data to better capture the relationships among modalities. Furthermore, we design the progressive alignment training strategy from primary to periphery.
3.2 Robust Audio Integration
The audio encoding component serves as a critical element in omni-modal models, as the system fundamentally relies on comprehending complex auditory inputs encompassing user speech signals and multimodal audio-visual comprehension data. In the Ola model, we implement a robust audio recognition module capable of processing speech, music, and video content, which seamlessly integrates with visual processing components.
Dual Audio Encoding. For audio encoding, we propose the dual encoder approach for the audio input. Specifically, we use Whisper-v3 [55] as the speech encoder and BEATs [8] as the music encoder for better alignment between audios and texts, providing richer audio information. The music encoder takes the original wav audio $A$ as inputs, while the speech encoder takes the wav transformed into Mel spectrogram representation $A_{(mel)}$ as inputs. Note that the Whisper encoder only supports a certain length of audio inputs, therefore we fix the sample rate as 16000Hz and cut the overlong audio into segments of 30 seconds $A_{1},A_{2},\cdots,A_{n}$ and conduct the encoder operation in batches $[f_{A_{1}},f_{A_{2}},\cdots,f_{A_{n}}]=\mathcal{E}_{A}([A_{1},A_{2},\cdots,A_{n}])$. The embedding features of the speech and music encoders are concatenated across channel dimensions for the comprehensive audio features $f_{A}$.
@@ 3.2 Robust Audio Integration
Mixing Audio Data. The audio training data are collected from various audio-relevant environments, including comprehensive speech and music understanding. For text-speech understanding, we design multiple tasks including ASR from LibriSpeech [51] and GigaSpeech [5] datasets, audio captioning from AudioCaps [29] and Clotho [14] datasets, speech question answering from LibriSpeech [51] datasets, audio question answering from WavCaps [46] and AudioCaps [29] datasets. For text-music understanding, we collect tasks including music captioning from MusicCaps [1], music question answering from MillionSong [45] and MusicNet [64]. The overall audio training data involves 1.1M samples. The relevant text question-answering representations are collected from SALMONN [61].
3.3 Bridging Modality Gaps for Omni-Modal Understanding